OpenStudy (anonymous):

Find dw/dt by converting w to a function of t and using the appropriate chain rule W = xycosz; x = t; y = t^2 ; z= arccost

OpenStudy (anonymous):

What is \(\cos(\cos^{-1}(x)))\)?

OpenStudy (anonymous):

Excuse the excess parenthesis.

OpenStudy (anonymous):

Question is incomplete. Are we supposed to find the dirivative or differential?

OpenStudy (anonymous):

No, \(\frac{dw}{dt}\) is the derivative.

OpenStudy (anonymous):

if you're talking about partial derivatives then the answer is 0

OpenStudy (anonymous):

...no.

OpenStudy (anonymous):

x'(t) then

OpenStudy (goformit100):

A warm Welcome to OpenStudy. I can help guide you through this useful site. You can ask your questions to me or you can message me. Please use the chat for off topic questions. Remember to give the person who helped you a medal by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it.